OpenStudy (anonymous):

Which of the following is the vertex form of the expression below? y = 9x2 - 90x + 231

OpenStudy (kira_yamato):

y=9(x^2-10x) + 231 y = 9(x - 5)^2 -45+231 y = 9(x-5)^2 + 186 Therefore vertex = (5,186)